# Volume effect on the extraction of sound velocity in high-energy nucleus-nucleus collisions

### v1

@@ IV Summary and discussion
With the help of the TRENTo modeling of initial states, we are able to investigate the effect of event-by-event fluctuations on the extraction of the sound velocity proposed in Ref. Gardim et al. (2020a). On one hand, these event-by-event fluctuations are crucial as they provide a means of variation of entropy in the systems created in ultra-central nucleus-nucleus collisions adiabatically. On the other hand, however, spatial inhomogeneity induced by these fluctuations results in variation of the effective volume. Accordingly, the extraction formula, which relates the speed of sound in QGP to the ratio of the relative change of mean transverse momentum and the relative change of the charged particle multiplicity, must receive corrections from the volume effect.
@@
Although the precise volume effect can only be calculated via model simulations, we propose an empirical formula (Eq. (11)) that can be used for realistic experiments, provided the mean transverse momentum and the charged particle multiplicity.

### v3

@@ IV Summary and discussion
With the help of the TRENTo modeling of initial states, we are able to investigate the effect of event-by-event fluctuations on the extraction of the speed of sound proposed in Ref. Gardim et al. (2020a). On one hand, these event-by-event fluctuations are crucial as they provide a means of variation of entropy in the systems created in ultra-central nucleus-nucleus collisions adiabatically. On the other hand, however, spatial inhomogeneity induced by these fluctuations results in variation of the effective volume. Accordingly, the extraction formula, which relates the speed of sound in QGP to the ratio of the relative change of mean transverse momentum and the relative change of the charged particle multiplicity, must receive corrections from the volume effect.